OpenStudy (anonymous):

write an equation of the line with the given slope and y intercept slope 3 and y-int (0,-3)

OpenStudy (katrinakaif):

Use the equation (y - y1) = m ( x-x1) as ( x1,y1) = ( 0,-3) and m = 3

OpenStudy (anonymous):

(y--3)=3(x-0)

OpenStudy (katrinakaif):

Now distribute the 3 to ( x - 0)

OpenStudy (anonymous):

how would i make the points to graph it ?

OpenStudy (katrinakaif):

y + 3 = 3x - 0 y = 3x + 3 To graph Make a point on the y-intercept 3, and use the slope 3/1 From the point 3 on the y-intercept, go 3 units up and one unit to the right and keep going 3 units up and one unit right until you make a long enough line.

OpenStudy (anonymous):

is it (-3,3)

OpenStudy (katrinakaif):

The equation is y = 3x + 3